# Polarization Patterns of Non-Circular Hotspots around Kerr Black Holes: A Preliminary Study

### v2

@@ 4.2 Results for hotspots from retrograde ISCO
Figure 6: The polarization image, the Q-U curve, and the net EVPA of the hotspot plunging from the retrograde ISCO, under the vertical magnetic field Eq. (3.18). The constant $\mathcal{B}$ within the Papapetrou-Wald potential is set to be 1. In the left two panels, the black hole spin is set to be $a=0.9M$. In the middle panel, we provide an enlarged view near the origin, where the scale of the coordinates should be multiplied by $10^{-10}$. In the right panel, the dashed line represents the projected retrograde ISCO on the observer’s screen, outside which $\chi_{net}$ is produced by the hotspots moving along retrograde circular orbits.
@@ 4.3 Results for homoclinic hotspots
Figure 7: The polarization images, the Q-U curves, and the net EVPAs of the hotspot traveling along a homoclinic orbit, under the vertical magnetic field (top row) and radial magnetic field (bottom row). In the left two columns, we set $a=0.9M,r_{c}=r_{ib}+0.15M$. For each case, the projected $r_{c}$ is denoted by the blue dashed circle on the screen; the blue segment of the Q-U curve corresponds to the portion of the homoclinic orbit moving outward from $r_{c}$ to $r_{3}$, while the red segment represents the portion of the orbit falling from $r_{3}$ to $r_{c}$. The right column presents the variation of the net EVPA of the hotspots falling from $r_{3}$ (the green point) to $r_{c}$, under different spins. For the ease of presentation, the radii of $r_{c}$ are chosen as $r_{c}=r_{ib}+1.2M/0.8M/0.15M$ for $a=0.1M/0.5M/0.9M$.
@@
Fig. 7 illustrates the polarization patterns of a hotspot along a homoclinic orbit, where the hotspot is perturbed to move outward from an unstable circular orbit. The net EVPA is greater under the radial magnetic field compared to the vertical magnetic field, a characteristic also observed in the case of an inspiraling hotspot. However, the Q-U curve exhibits a distinctive multi-loop structure, as illustrated in the middle column of Fig. 7. In the previous studies, the Q-U loop has been considered to be a significant characteristic in the hotspot astrometry, attributed to the nonzero observation angle [13]. We now highlight that even when being observed face-on, such a loop structure can arise from a homoclinic orbit. It is crucial to note that one-loop structure in the Q-U curves is a general feature of elliptical-like orbits, depending solely on the variation of total intensity along the orbit and the completion of a full period in $\varphi$. However, only the homoclinic orbit can generate a multi-loop structure on the Q-U plane, where the hotspot wraps around $r_{c}$ an infinite number of times. Furthermore, the black hole spin introduces a slight breaking of mirror symmetry in the homoclinic orbit [41]. This asymmetry is amplified in the corresponding Q-U curve, as can be seen in the right column of Fig. 7.
@@ 4.4 Results for accretion flows within the plunging region
In the intra-ISCO area of a thin accretion disk, the fluid inevitably adopts a plunging trajectory towards the black hole’s horizon, forming a distinct plunging region. This region holds significant astrophysical implications [52, 39, 42]. Notably, it plays a crucial role in explaining the observational spectrum of accretion disks in the $X$-ray band [38, 40]. If the plunging region is present near a supermassive black hole like M87* or SgrA*, it can also yield unique optical signatures in the radio band, potentially detectable by the EHT. Within this plunging region, the fluid dynamics are primarily governed by the gravitational pull, causing the streamlines to closely follow timelike geodesics. The analytical framework outlined in Sec.2 and Sec. 3 can be directly applied to study the plunging region. For example, superimposing the trjectories of Eq. (2.2) with varying initial azimuthal angles can depict a stable, axisymmetric profile of an intra-ISCO accretion disk.
Figure 8: The observed total intensity and the net EVPA for a plunging-region disk comprised of entirely zero-conductivity plasma. We set the radial velocity at the ISCO crossing as $\delta{U}=10^{-3}$. For each case, the dashed line denotes the location of the inner shadow calculated using Eq. (3.10).
In a typical accretion disk, the magnetic field is intricately linked with the fluid dynamics. In the case of a plasma with zero conductivity, the accompanying magnetic field is governed by the ideal-MHD condition [50], resulting in
@@
Here $g$ is the determinant of the metric. Note that $B^{r},B^{\phi}$ diverge at the ISCO where $U^{r}=0$. This divergence should be addressed by introducing a non-zero radial velocity when crossing the ISCO, as suggested by [42], given by $U^{r}=U^{r}_{I}-\delta U$, where $\delta U\ll 1$ is a small correction based on Eq. (2.2). In the zero-conductivity fluid, the magnetic field lines align with the streamlines, leading to a negligible vertical component near the equatorial plane. However, real accretion flows may exhibit some conductivity near the equatorial plane, resulting in a non-zero vertical component there. For simplicity, the Papapetrou-Wald potential given by Eq. (3.18) can be utilized to represent a typical vertical magnetic field within the plunging region. Consequently, the results depicted in the top row of Fig. 5 can be interpreted as the polarization pattern of the plunging region with a vertical magnetic field. The suppression of central intensity outlines the horizon on the screen, commonly referred to as the inner shadow of the accretion disk [37].
@@
Under the accompanying magnetic field of Eq. (4.3), the optical appearance of the disk is illustrated in Fig. 8, where we have chosen $\delta{U}=10^{-3}$ as a conservative estimation. The behavior of $\mathcal{I}_{o}$ as the radius varies reveals a significant reduction in the observed intensity not so close to the horizon. For instance, $\mathcal{I}_{o}$ at $r=(r_{+}+r_{I})/2$ diminishes to nearly $1/10000$ of the value at $r=r_{I}$. Thus, with such a magnetic field configuration, the plunging region appears much darker, resulting in a larger central dark region on the screen compared to the inner shadow produced by the vertical field. Regarding the polarization pattern, as the radius decreases from the ISCO, $\chi_{net}$ remains less than $\pi/2$ across all scenarios under the vertical magnetic field, as depicted in the top-right panel of Fig. 5, indicating a counterclockwise pattern. Conversely, with the accompanying magnetic field, $\chi_{net}$ is close to 00 or $\pi$, indicating a highly radial polarization pattern, as shown in the right panel of. Fig. 8. Furthermore, for $a=0.1M$ or $0.5M$, the net EVPA undergoes a transition from weakly clockwise to weakly counterclockwise patterns, manifesting as jumps from $180^{\circ}$ to $0^{\circ}$. For $a=0.9M$ the net EVPA exhibits a weakly clockwise pattern.
@@ 5 Summary
In this study, we studied the polarized images of luminous hotspots in close proximity to the horizon of a Kerr black hole. We focused on the thermal synchrotron emission from the hotspots moving along the plunging orbits and the homoclinic orbits [44, 41], while considering two distinct magnetic field configurations: the vertical and radial fields [48, 50]. For an on-axis observer, we introduced a new approximate function to facilitate the computation of the light bending, thereby simplifying the calculation of the imaging process. To elucidate the polarization patterns, we conducted the calculations involving the Stokes parameters and derived the total intensity and the polarization vector along a hotspot’s trajectory. The main results of this study are listed below:
  • During the inspiral of a hotspot from the prograde ISCO, our study revealed that when subject to a vertical magnetic field, the total intensity of the hotspot decreases monotonically as it approaches the horizon, while the net EVPA continually increases but remains $\chi_{net}<\pi/2$. In contrast, with a radial magnetic field, the total intensity does not follow monotonic patterns, and the net EVPA always satisfies $\chi_{net}>\pi/2$. The Stokes parameters demonstrate a right-handed spiral shape on the Q-U plane, with the phase determined by the hotspot’s motion in the $\phi$ direction.
@@
Moreover, by considering the plunging region of an accretion flow approximated by plunging geodesics, we extended our study to the polarized images of the intra-ISCO region of a thin disk. Our study provided, for the first time, the details of the polarizations of a plunging region disk under different magnetic field configurations. Under the vertical magnetic field or the accompanying magnetic field given by Eq. (4.3), the optical appearances of the disk are totally different, as illustrated in Fig. 5 and Fig. 8. Upon comparing the total intensities between the two scenarios, we observed that the plunging region with the accompanying magnetic field appears notably darker than when under the vertical magnetic field. Furthermore, the net EVPA under the vertical magnetic field indicates a counterclockwise pattern, whereas the net EVPA under the accompanying magnetic field exhibits a highly radial pattern.
We end this work by outlining several limitations and future perspectives of this study. Firstly, for the purpose of an analytical study, we considered the hotspots traveling along the plunging and the homoclinic geodesics with integrable trajectory functions. It is essential to consider the hotspots with other types of motion, such as the bound orbits and the deflecting orbits [53], or the hotspots not confined to the equatorial plane. Secondly, although the inclination angles for observing M87* and Sgr A* are small, it is still necessary to consider the effects of a nonzero inclination, especially how it distorts the geometric shape of the Q-U curve [13]. Lastly, we neglected the multi-level images of the hotspots or the inner-ISCO disk due to strong gravitational lensing in this study, which manifest as highly distorted curves closely following the critical curve of the Kerr black hole. The lensed images could contribute both to the total flux and the observed Stokes parameters.

### v3

@@ 4.2 Results for hotspots from retrograde ISCO
Figure 6: The time-integrated polarized image, the Q-U curve, and the net EVPA of the hotspot plunging from the retrograde ISCO, under the vertical magnetic field Eq. (3.18). The constant $\mathcal{B}$ within the Papapetrou-Wald potential is set to be 1. In the left two panels, the black hole spin is set to be $a=0.9M$. In the middle panel, we provide an enlarged view near the origin, where the scale of the coordinates should be multiplied by $10^{-10}$. In the right panel, the dashed line represents the projected retrograde ISCO on the observer’s screen, outside which $\chi_{net}$ is produced by the hotspots moving along retrograde circular orbits.
@@ 4.3 Results for homoclinic hotspots
Figure 7: The time-integrated polarized images, the Q-U curves, and the net EVPAs of the hotspot traveling along a homoclinic orbit, under the vertical magnetic field (top row) and radial magnetic field (bottom row). In the left two columns, we set $a=0.9M,r_{c}=r_{ib}+0.15M$. For each case, the projected $r_{c}$ is denoted by the blue dashed circle on the screen; the blue segment of the Q-U curve corresponds to the portion of the homoclinic orbit moving outward from $r_{c}$ to $r_{3}$, while the red segment represents the portion of the orbit falling from $r_{3}$ to $r_{c}$. The right column presents the variation of the net EVPA of the hotspots falling from $r_{3}$ (the green point) to $r_{c}$, under different spins. For the ease of presentation, the radii of $r_{c}$ are chosen as $r_{c}=r_{ib}+1.2M/0.8M/0.15M$ for $a=0.1M/0.5M/0.9M$.
@@
Fig. 7 illustrates the polarization patterns of a hotspot along a homoclinic orbit, where the hotspot is perturbed to move outward from an unstable circular orbit. The net EVPA is greater under the radial magnetic field compared to the vertical magnetic field, a characteristic also observed in the case of an inspiraling hotspot. However, the Q-U curve exhibits a distinctive multi-loop structure, as illustrated in the middle column of Fig. 7. In the previous studies, the Q-U loop has been considered to be a significant characteristic in the hotspot astrometry, attributed to the nonzero observation angle [13]. We now highlight that even when being observed on-axis, such a loop structure can arise from a homoclinic orbit.
It is crucial to note that one-loop structure in the Q-U curves is a general feature of elliptical-like orbits, depending solely on the variation of total intensity along the orbit and the completion of a full period in $\varphi$. However, only the homoclinic orbit can generate the multi-loop structure on the Q-U plane in Fig. 7, where the hotspot wraps around $r_{c}$ infinite times. Furthermore, the black-hole spinning introduces a slight breaking of mirror symmetry in the homoclinic orbit [43]. This asymmetry is amplified in the corresponding Q-U curve, as can be seen in the right column of Fig. 7. Also, we want to stress that the hotspot itself may undergo evolution, causing variations in its emission profile. This may lead to the Q-U pattern evolving between subsequent orbits of the hotspot on a circular orbit. The difference in the loop structures in the Q-U plane caused by the homoclinic orbit and by the evolution of the hotspot is an interesting issue worthy of further investigation.
4.4 Effect of non-zero inclination angle
In this section, we give a preliminary study on the effect of non-zero inclination angle on the polarization pattern of a hotspot. Since the imaging formula in Sec.3.2 is not applicable for inclined observers, we employ the numerical backward ray-tracing method, as described in [14, 20]. Besides, we use the Cartesian coordinates $(\alpha,\beta)$ introduced in [57] to cover the screen.
Figure 8: The polarization structures of hotspots along a prograde plunging orbit under a vertical magnetic field. From left to right, the black hole spin is set to $a=0.1M,0.5M,$ and $0.9M$. From top to bottom, the inclination angle is set to $17^{\circ},30^{\circ},$ and $60^{\circ}$.
In Figure 8, we illustrate the polarization vectors along the projected trajectory of a plunging hotspot on the screen, assuming a vertical magnetic field. From left to right columns, the black hole spin is set to be $0.1M,0.5M,$ and $0.9M$. From top to bottom rows, the inclination angle is set to be $17^{\circ},30^{\circ},$ and $60^{\circ}$. As the inclination angle increases, the projection of the ISCO on the screen gradually transforms from a circle to an elliptical shape, leading to the corresponding changes in the appearance of the hotspot’s image. It is worth noting that for small inclination angles, such as $\theta=17^{\circ}$ and $30^{\circ}$, the results exhibit only minor deviations from the one in the on-axis case. Given that current observations of M87 and Sgr A* suggest a preference for small inclination angles, our previous discussion for the on-axis case remains qualitatively valid.
Figure 9: The observed total intensity and net EVPA along a prograde plunging orbit under a vertical magnetic field. From top to bottom, the inclination angle is set to $17^{\circ},30^{\circ},$ and $60^{\circ}$. The maximum value of the total intensity has been normalized to 1. The maximum value of $\varphi$ corresponds to $r_{s}=r_{+}+0.05M$ and the minimum value of $\varphi$ corresponds to $r_{s}=r_{ms}-0.1M$.
In Figure 9, we illustrate the variations in the observed total intensity and net EVPA along the projected trajectory of the hotspot. To describe the projected trajectory, we introduce polar coordinates defined as $\rho=\sqrt{\alpha^{2}+\beta^{2}},\varphi=\arctan{(\beta/\alpha)}$. For nonzero inclination angles, the angular radius $\rho$ along the projected trajectory on the observer’s screen is no longer a monotonic parameter, in contrast to the on-axis case. Consequently, we plot the observed quantities as the functions of the polar angle $\varphi$ on the screen. As the hotspot inspirals counterclockwise toward the black hole, $\varphi$ increases. A clear oscillatory behavior in both $I_{o}$ and $\chi_{net}$ with respect to $\varphi$ is evident, which can be attributed to the azimuthal angle $\phi$ completing multiple full $360^{\circ}$ rotations around the black hole.
In Figure 9, each peak in the oscillation of $I_{o}$ corresponds to a position near the left side of the $\alpha$-axis on the observer’s screen, resulting from the Doppler effect as the hotspot moves toward the observer. This effect becomes more pronounced with increasing inclination angle. Additionally, the variations in the observed quantities across different oscillation cycles reflect the hotspot’s gradual deviation from the ISCO. As the hotspot plunges into the black hole, the heights of the intensity peaks progressively decrease, a trend that is more pronounced in high-spin cases. Regarding the polarization pattern, it generally retains a counterclockwise orientation. As the inclination angle increases, the oscillation amplitude of $\chi_{net}$ also grows. For an inclination of $60^{\circ}$, a portion of the polarization in the upper half of the screen exhibits a clockwise pattern. Moreover, the variation in $\chi_{net}$ is smoothest when $a=0.9M$, regardless of the inclination angle.
@@ 4.5 Results for accretion flows within the plunging region
In the intra-ISCO area of a thin accretion disk, the fluid inevitably adopts a plunging trajectory towards the black hole’s horizon, forming a distinct plunging region. This region holds significant astrophysical implications [58, 41, 44]. Notably, it plays a crucial role in explaining the observational spectrum of accretion disks in the X-ray band [40, 42]. If the plunging region is present near a supermassive black hole like M87* or SgrA*, it can also yield unique signatures in the radio band, potentially detectable by the EHT. Within this plunging region, the fluid dynamics are primarily governed by the gravitational pull, causing the streamlines to closely follow timelike geodesics. The analytical framework outlined in Sec.2 and Sec. 3 can be directly applied to study the plunging region. For example, superimposing the trjectories of Eq. (2.2) with varying initial azimuthal angles can depict a stable, axisymmetric profile of an intra-ISCO accretion disk.
Figure 10: The observed total intensity and the net EVPA for a plunging-region disk comprised of entirely zero-conductivity plasma. We set the radial velocity at the ISCO crossing as $\delta{U}=10^{-3}$. For each case, the dashed line denotes the projection of the event horizon calculated using Eq. (3.10).
In a typical accretion disk, the magnetic field is intricately linked with the fluid dynamics. In the case of a plasma with infinite conductivity, the accompanying magnetic field is governed by the ideal-MHD condition [55], resulting in
@@
Here $g$ is the determinant of the metric. Note that $B^{r},B^{\phi}$ diverge at the ISCO where $U^{r}=0$. This divergence should be addressed by introducing a non-zero radial velocity when crossing the ISCO, as suggested by [44], given by $U^{r}=U^{r}_{I}-\delta U$, where $\delta U\ll 1$ is a small correction based on Eq. (2.2). In the zero-conductivity fluid, the magnetic field lines align with the streamlines, leading to a negligible vertical component near the equatorial plane. However, real accretion flows may exhibit some conductivity near the equatorial plane, resulting in a non-zero vertical component there. For simplicity, the Papapetrou-Wald potential given by Eq. (3.18) can be utilized to represent a typical vertical magnetic field within the plunging region. Consequently, the results depicted in the top row of Fig. 5 can be interpreted as the polarization pattern of the plunging region with a vertical magnetic field. The suppression of central intensity outlines the horizon on the screen, commonly referred to as the inner shadow of the accretion disk [59, 39].
@@
Under the accompanying magnetic field of Eq. (4.3), the image of the disk seen by an on-axis observer is illustrated in Fig. 10, where we have chosen $\delta{U}=10^{-3}$ as a conservative estimation. The behavior of $\mathcal{I}_{o}$ as the radius varies reveals a significant reduction in the observed intensity not so close to the horizon. For instance, $\mathcal{I}_{o}$ at $r=(r_{+}+r_{I})/2$ diminishes to nearly $1/10000$ of the value at $r=r_{I}$. Thus, with such a magnetic field configuration, the plunging region appears much darker, resulting in a larger central dark region on the screen compared to the inner shadow produced by the vertical field. Regarding the polarization pattern, as the radius decreases from the ISCO, $\chi_{net}$ remains less than $\pi/2$ across all scenarios under the vertical magnetic field, as depicted in the top-right panel of Fig. 5, indicating a counterclockwise pattern. Conversely, with the accompanying magnetic field, $\chi_{net}$ is close to 00 or $\pi$, indicating a highly radial polarization pattern, as shown in the right panel of. Fig. 10. Furthermore, for $a=0.1M$ or $0.5M$, the net EVPA undergoes a transition from weakly clockwise to weakly counterclockwise patterns, manifesting as jumps from $180^{\circ}$ to $0^{\circ}$. For $a=0.9M$ the net EVPA exhibits a weakly clockwise pattern.
4.6 Higher-order images
In the preceding discussion, we did not consider the higher-order images. Here, we briefly explain why we neglect these lensed images. The higher-order images arise from the photons that undergo multiple orbits around the black hole before reaching the observer. These images are confined within the so-called lensing band, whose inner and outer boundaries correspond to the lensed images of circles at $r=r_{+}$ and $r\rightarrow\infty$ on the equatorial plane respectively[51, 60]. For small inclination angles, the lensing band is located very close to the critical curve, which is characterized by the photon-sphere radius [61]:
$\tilde{r}_{0}=M+2\sqrt{M^{2}-\frac{a^{2}}{3}}\cos{\mathopen{}\mathclose{{}\left[\frac{1}{3}\arccos{\frac{M\mathopen{}\mathclose{{}\left(M^{2}-a^{2}}\right)}{\mathopen{}\mathclose{{}\left(M^{2}-\frac{a^{2}}{3}}\right)^{3/2}}}}\right]}\,,$
$\tilde{\rho}_{0}=\sqrt{\frac{\tilde{r}_{0}^{3}}{a^{2}}\mathopen{}\mathclose{{}\left[\frac{4M(\tilde{r}_{0}-r_{+})(\tilde{r}_{0}-r_{-})}{(\tilde{r}_{0}-M)^{2}}-\tilde{r}_{0}}\right]+a^{2}}\,.$ (4.4)
As the spin decreases from $a=M$ to 00, $\tilde{r}_{0}$ increases from $2.414M$ to $3M$. The radius of the critical curve, given by $\rho=\tilde{\rho}_{0}$, is determined by the critical impact parameters and varies from $4.828M$ to $3\sqrt{3}M$.
For the direct image, the angular radius follows the approximate relation $\rho\approx r/M+1$ (see Eq. (3.11)) for small inclination angles, indicating that the direct image experiences minimal distortion. As the plunging orbit extends from the ISCO to the event horizon, the inner and outer boundaries of the time-integrated direct image are given by $\rho_{d}(r_{+})\approx r_{+}/M+1$, $\rho_{d}(r_{ms})\approx r_{ms}/M+1$, where and hereafter we use $\rho_{d}(r)$, $\rho_{l}(r)$ to denote the angular radii of the direct and lensed images, respectively, on the observer’s screen. As the spin decreases from $a=M$ to 00, $\rho_{d}(r_{+})$ ranges from $2M$ to $3M$, while $\rho_{d}(r_{ms})$ varies from $2M$ to $7M$. Consequently, as the spin decreases, $\rho_{d}(r_{ms})$ decreases faster than $\tilde{\rho}_{0}$. If the spin is sufficiently high such that $\rho_{d}(r_{ms})<\tilde{\rho}_{0}\approx\rho_{l}(r)$, then the time-integrated direct and higher-order images remain spatially separated on the screen. A precise calculation shows that the critical case occurs at $a=0.5729M$, where we have $\rho_{d}(r_{ms})=\rho_{l}(r_{+})$. Therefore, as long as $a>0.5729M$, the time-integrated direct and higher-order images can be analyzed independently, as illustrated in Figure 11.
Figure 11: The direct images and the higher-order images. The blue curves represent the projections of the ISCO on the observer’s screen, while the red curves correspond to the projections of the event horizon. Solid and dashed curves indicate their direct and lensed images, respectively. Besides, we shaded the direct image region in light blue and the lensed image region in light yellow. From left to right, the black hole spin is set to $a=0.1M,0.5729M,0.9M$. The rightmost panel illustrates the variations of $\rho_{d}(r_{+}),\rho_{d}(r_{ms}),\rho_{l}(r_{+}),\rho_{l}(r_{ms})$ as the functions of spin. The lensed images are computed using Eq. (78) in [51]. When $a>0.5729M$, we have $\rho_{d}(r_{+})<\rho_{d}(r_{ms})<\rho_{l}(r_{+})<\rho_{l}(r_{ms})$, which means the lensed image is separated from the direct image.
Furthermore, we note that the emission from the hotspot is inherently a dynamical process. In astrophysics, the typical timescale for plunging objects in Kerr black holes is approximately $t_{\text{pl}}=2M\sim 5M$ [44]. In contrast, the time delay between the photons forming adjacent-order images is given by $t_{\text{dl}}\approx 3\sqrt{3}\pi M$ [51]. Since $t_{\text{dl}}>t_{\text{pl}}$, the direct image of the plunging hotspot will be observed before the lensed photons for higher-order images reach the observer. Consequently, the direct and higher-order images of the plunging hotspot remain separated in the time domain.
@@ 5 Summary
We studied the polarized images of luminous hotspots on non-circluar orbits, in close proximity to the horizon of a Kerr black hole. We focused on the synchrotron emission from the hotspots moving along the plunging orbits and the homoclinic orbits [46, 43], while considering two distinct magnetic field configurations: the vertical and radial fields [53, 55]. For an on-axis observer, we introduced a new approximate function to facilitate the computation of the light bending, thereby simplifying the calculation of the imaging process. To elucidate the polarization patterns, we conducted the calculations involving the Stokes parameters and derived the total intensity and the polarization vector along a hotspot’s trajectory. The main results of our study are listed below:
  • During the inspiral of a hotspot from the prograde ISCO, when subject to a vertical magnetic field, the total intensity of the hotspot decreases monotonically as it approaches the horizon, while the net EVPA continually increases but remains $\chi_{net}<\pi/2$. In contrast, with a radial magnetic field, the total intensity does not follow monotonic patterns, and the net EVPA always satisfies $\chi_{net}>\pi/2$. The Stokes parameters demonstrate a right-handed spiral shape on the Q-U plane, with the phase determined by the hotspot’s motion in the $\phi$ direction.
@@
  • For a nonzero inclination angle, the time-integrated image takes on an elliptical shape, with the total intensity significantly influenced by the Doppler effect. The polarization pattern is generally counterclockwise; however, for large inclination angles ($\gtrsim 60^{\circ}$), a portion of the net EVPA in the upper half of the screen exhibits a clockwise pattern. Additionally, a higher spin more effectively smooths out the variations in the net EVPA along the trajectory.
Moreover, by considering the plunging region of an accretion flow approximated by plunging geodesics, we extended our study to the polarized images of the intra-ISCO region of a thin disk. Our study provided, for the first time, the details of the polarizations of a plunging region disk under different magnetic field configurations. Under the vertical magnetic field or the accompanying magnetic field given by Eq. (4.3), the images of the disk are totally different, as illustrated in Fig. 5 and Fig. 10. Upon comparing the total intensities between the two scenarios, we observed that the plunging region with the accompanying magnetic field appears notably darker than when under the vertical magnetic field. Furthermore, the net EVPA under the vertical magnetic field indicates a counterclockwise pattern, whereas the net EVPA under the accompanying magnetic field exhibits a highly radial pattern.
We end this work by outlining several limitations and future perspectives of present study. In order to carry an analytical study, we only considered the hotspots traveling along the plunging and the homoclinic geodesics with integrable trajectory functions. It is essential to consider the hotspots with other types of motions, such as the bound orbits and the deflecting orbits [62], or the hotspots not confined to the equatorial plane. Additionally, we have neglected the effects of higher-order images. In the high-spin cases, the time-integrated higher-order images are well separated from the direct one and do not significantly affect the results. However, in the low-spin cases, as there are overlap between the time-integrated direct and higher-order images, one must take the higher-order images seriously. Moreover, the higher-order images may contribute to the polarized light curve as well, and they should be treated with care. We would like to leave these issues to future studies.